Waterways Continuing Problem 11 a-g

Waterways Corporation uses very stringent standard costs in evaluating its manufacturing efficiency. These standards are not “ideal” at this point, but the management is working toward that as a goal. At present, the company uses the following standards.
Materials
Item Per unit Cost
Metal 1 lb. 63¢ per lb.
Plastic 12 oz. $1.00 per lb.
Rubber 4 oz. 88¢ per lb.
Direct labor
Item Per unit Cost
Labor 15 min. $9.00 per hr.
Predetermined overhead rate based on direct labor hours = $3.62

The January figures for purchasing, production, and labor are:
The company purchased 218,100 pounds of raw materials in January at a cost of 79¢ a pound.
Production used 218,100 pounds of raw materials to make 110,000 units in January.
Direct labor spent 18 minutes on each product at a cost of $8.80 per hour.
Overhead costs for January totaled $30,265 variable and $74,000 fixed.

Answer the following questions about standard costs.
What is the materials price variance? (Round per unit calculations to 2 decimal places, e.g. 1.25 and final answer to 0 decimal places, e.g. 125.)
Materials price variance = ? $

FavorableUnfavorableNeither favorable nor unfavorable

What is the total materials variance? (Round per unit calculations to 2 decimal places, e.g. 1.25 and final answer to 0 decimal places, e.g. 125.)
Total materials variance = ? $

Neither favorable nor unfavorableUnfavorableFavorable

What is the labor price variance? (Round per unit calculations to 2 decimal places, e.g. 1.25 and final answer to 0 decimal places, e.g. 125.)
Labor price variance = ? $

FavorableUnfavorableNeither favorable nor unfavorable

What is the labor quantity variance? (Round per unit calculations to 2 decimal places, e.g. 1.25 and final answer to 0 decimal places, e.g. 125.)
Labor quantity variance = ? $

UnfavorableNeither favorable nor unfavorableFavorable

What is the total labor variance? (Round per unit calculations to 2 decimal places, e.g. 1.25 and final answer to 0 decimal places, e.g. 125.)
Total labor variance = ? $

UnfavorableNeither favorable nor unfavorableFavorable

What is the total overhead variance? (Round per unit calculations to 2 decimal places, e.g. 1.25 and final answer to 0 decimal places, e.g. 125.)
Total overhead variance = ?
Answer #1

Solution

Material Standards set for 1 unit,

Metal = 1 lb. @ 63 cent, or $ 0.63 per lb. = $ 0.63

Plastic = 12 oz., or 0.75 lb. @ $ 1 per lb. = $ 0.75

Rubber = 4 oz., or 0.25 lb. @ 88 cent, or $ 0.88 per lb = $ 0.22

Therefore, Standard Quantity for 1 unit = (1 + 0.75 + 0.25) = 2 lbs.

Standard Rate = $ (0.63 + 0.75 + 0.22) = $ 1.60

The company bought 218100 lbs. of Materials @ 79 cent, or $ 0.79 per lb. to make 110000 units

Therefore, Standard Quantity to be used (SQ) = 2 lbs. X 110000 units = 220000 lbs.

Actual Quantity used (AQ) = 218100 lbs.

Standard Rate per lb. (SR) = $ 1.60

Actual Rate per lb. (AR) = $ 0.79

Now,

(i) Material Price Variance = (SR - AR) X AQ

= $ (1.60 - 0.79) X 218100 = $ 176661 (F)

(ii) Total Material Variance = (SR X SQ) - (AR X AQ)

= $ [ (1.60 X 220000) - (218100 X 0.79) ]

= $ ( 352000 - 172299 )

= $ 179701 (F)

Labour Standards Set,

Standard Hour (SH) = 15 Mins., or 0.25 Hour per unit produced, or 0.25 X 110000 = 27500 Hours

Standard Rate (SR) = $ 9

Actual Hour (AH) = 18 Mins., or 0.30 Hour per unit, or 0.30 X 110000 = 33000 Hours

Actual Rate (AR) = $ 8.80

Now,

(iii) Labour Price Variance = (SR-AR)XAH

= $ (9 - 8.8) X 33000

= $ 6600 (F)

(iv) Labour Quantity Variance = (SH - AH) X SR

= $ (27500 - 33000) X 9

= $ 49500 (U)

(v) Total Labour Variance = (SH X SR) - (AH X AR)

= $ [ (27500 X 9) - (33000 X 8.80) ]

= $ (247500 - 290400)

= $ 42900 (U)

For Overheads,

Standard Set = $ 3.62 per actual direct labour hour spent, or 33000 hours = 33000 X $ 3.62 = $ 119460

Overhead Absorbed = $ (30265 + 74000) = $ 104265

Therefore,

(vi) Total Overhead variance = Standard Overhead - Actual Overhead

= $ (119460 - 104265)

= $ 15195 (F)